# OBI-Bench: Can LMMs Aid in Study of Ancient Script on Oracle Bones?

### v1

@@ Abstract
We introduce OBI-Bench, a holistic benchmark crafted to systematically evaluate large multi-modal models (LMMs) on whole-process oracle bone inscriptions (OBI) processing tasks demanding expert-level domain knowledge and deliberate cognition. OBI-Bench includes 5,523 meticulously collected diverse-sourced images, covering five key domain problems: recognition, rejoining, classification, retrieval, and deciphering. These images span centuries of archaeological findings and years of research by front-line scholars, comprising multi-stage font appearances from excavation to synthesis, such as original oracle bone, inked rubbings, oracle bone fragments, cropped single character, and handprinted character. Unlike existing benchmarks, OBI-Bench focuses on advanced visual perception and reasoning with OBI-specific knowledge, challenging LMMs to perform tasks akin to those faced by experts. The evaluation of 6 proprietary LMMs as well as 17 open-source LMMs highlights the substantial challenges and demands posed by OBI-Bench. Even the latest versions of GPT-4o, Gemini 1.5 Pro, and Qwen-VL-Max are still far from public-level humans in some fine-grained perception tasks. However, they perform at a level comparable to untrained humans in deciphering task, indicating remarkable capabilities in offering new interpretative perspectives and generating creative guesses. We hope OBI-Bench can facilitate the community to develop domain-specific multi-modal foundation models towards ancient language research and delve deeper to discover and enhance these untapped potentials of LMMs.
@@ 1 Introduction
Nowadays, the emergent large multi-modal models (LMMs) have brought opportunities for solving multidisciplinary tasks with powerful visual perception, understanding, and reasoning abilities (Yue et al., 2024; Caffagni et al., 2024; OpenAI, 2024b). Meanwhile, the ability of LMMs remains unclear on fine-grained perception and cognition (i.e., high-level perception), which play significant roles in interpreting ancient texts and its associated tasks on image material pre-processing. The perception at different granularities is strongly associated with a wide range of OBI tasks, such as distinguishing noise from real oracle bone character recognition (Wang et al., 2022), marginal sealability checking in rejoining (Zhang et al., 2021c) and component-level structural decomposition in retrieval (Hu et al., 2024). Henceforth, it is worth evaluating the current abilities of LMMs in fine-grained perception and cognition based on actual needs to relieve extensive human resources and seek feasible solutions for future OBI research.
@@
To systematically evaluate the whole-process perception ability on various visual granularity under diverse OBI tasks, we collect 5,523 OBI images from 11 distinct sources. Due to the lack of publicly available OBI recognition datasets on real oracle bones and OBI rejoining datasets, we propose the original oracle bone recognition (O2BR) dataset and OBI-rejoin dataset as complements. Aligned with existing practices (Yue et al., 2024; Wu et al., 2024), each image in OBI-Bench is equipped with a question alongside a correct answer. Moreover, we craft a spectrum of questions: What question, Yes-or-No question, How question, and Where question, evaluating from coarse-grained perception to finer-grained perception. A total of 23 popular LMMs are selected for evaluation including 6 proprietary LMMs (e.g., GPT-4o, Qwen-VL-Max, and Gemini 1.5 Pro) and 17 open-source LMMs (e.g., LLaVA-Next, InternVL2, and mPLUG-Owl3). We showcase some empirical findings here: 1) LMMs still have much room for improvement in fine-grained OBI recognition tasks such as quantity detection or locating; 2) LMMs are not sensitive enough to local information such as the borders of the fragmented margin to directly meet the requirements of rejoining. Meanwhile, part of LMMs can help OBI scholars narrow down the range of qualified fragments (GPT-4o and Gemini 1.5 Pro achieves an average performance of 77.57% on the Acc@10 metric); 3) LMMs can be applied in classification and retrieval tasks, especially for cleaner handprinted datasets; 4) partial LLMs rival the performance of public-level humans in deciphering tasks and exhibit remarkable visual reasoning ability in interpreting those genuine undeciphered oracle bone characters. The full list of our findings from different OBI tasks is in the Sec. 3.
@@ 2.1 General Principles
Covering Multi-stage Font Appearances. To cover diverse appearances of the oracle bones since their excavation, we collect multi-sourced images for each task, as depicted in Tab. 1. In particular, considering that there are currently no available original oracle bone recognition datasets and rejoining datasets, we proposed the O2BR and OBI-rejoin datasets respectively (See more details in App. B). Due to the specific requirement of domain knowledge in OBI research, three domain experts and one senior OBI scholar were involved in the annotation process of O2BR and OBI-rejoin as well as the sampling process of all source datasets. We include all the manifestations of the entire process of OBI processing from excavation (i.e., original oracle bone) to artificial synthesis (i.e., handprinted or computer-generated). The diverse and multiple sources of images morph the OBI-Bench into a holistic and balanced benchmark to fairly evaluate the competency of LMMs in OBI-related tasks.
@@ 3.1 Evaluation on Recognition
Setup. To evaluate how well LMMs perform under various OBI recognition scenarios, we design four evaluation schemes across two image sources (i.e., original oracle bone from our self-curated O2BR dataset and inked rubbings from the YinQiWenYuandetection (AYNU, 2020) dataset): (1) evaluating via the What query to measure the preciseness of the global perceptual descriptions on OBIs. Given the commonality of contents in sampled test sets, we predefined reference descriptions for each of the two datasets. The max_tokens for all LMMs is set to $200$. The similarity between text embeddings is used as the evaluation metric. (2) evaluating via the Yes-or-No query after specifying the response direction (the words “OBI” appear explicitly in the query), designed to complement the evaluation on coarse-grained perception. The accuracy of answers is adopted as the evaluation criterion. (3) evaluating via the How query to quantify the ability of LMMs in parsing irregularly structured character-level OBI contents. We use mean relative error (MRE) as the metric. (4) evaluating via the Where query to achieve finer-grained perception (i.e., dense oracle bone character locating), designed to more effectively uncover the practicability of LMMs in real OBI processing pipeline. Mean intersection over union (mIoU) is used as the metric. More details are in App. C.2.
@@
Results. In scenario (1), as shown in Tab. 2, among proprietary LMMs, the recently-released GPT-4o reaches the best performance in terms of the relevance of answers on coarse-grained perception, followed by Qwen-VL-Max, which shows rather close results. By achieving over 12.4% on average performance improvement, these models provide more precise answers than the open-source LMMs. Similar improvement can be observed for the same series of LMM (e.g., LLaVA and InternVL2), as the number of parameters of the language backbone increases. In scenario (2), we add keyword (i.e., oracle bone) directly to a Yes-or-No question to further evaluate the global perception of LMMs. We find most LMMs perform nearly 100% correctness, indicating that these models are highly effective in handling queries that involve explicitly directed content. One exception is the GLM-4V series that only achieves an average accuracy of 34.34%. We speculate it is due to the answer preference bias (Wu et al., 2024). Deeper analysis is in App. C.2.2. In scenario (3), we transition to examining the fine-grained visual perceptual of LMMs, i.e., counting the number of oracle bone characters. Our results in Tab. 2 indicate that GLM-4V and InternVL2-Llama3-76B exhibit the best performance within their respective division groups. Additionally, we observe some abnormally large values (>50) or repeated values from Gemini 1.5 Flash, xGen-MM-4B, LLaVA-NeXT-72B, LLaVA-v1.5-13B, and Qwen-VL that results in their relatively large MRE (Zhou et al., 2024). In scenario (4), all evaluated LMMs consistently underperform. Gemini 1.5 Pro significantly outperforms other LMMs (surpasses GPT-4v and GPT-4o two orders of magnitude) (Ueno & Lynn, 2024) with its exclusive coordinate return function. Overall, current LMMs are still not usable for character-level OBI locating and are far from the public-level human.
@@ 3.2 Evaluation on Rejoining
Setup. Given the absence of publicly available rejoining datasets in OBI domain, we create a dataset consisting of 200 complete OBI pieces across two appearances (i.e., original and inked oracle bone fragments) with 483 rejoinable fragments. More details are in App. B.2. To evaluate the rejoining performance of LMMs on oracle bone fragments with different marginal sealability concerns, we elaborate two evaluation scenarios: (1) evaluating via standard Yes-or-No question that represents the absolute judgments to tentatively explore the rejoining performance; (2) evaluating via How question that outputs rejoinable probabilities for pairs of fragments, allowing for a more accurate and quantitative measurement. Accuracy@k indices are utilized as evaluation metrics.
@@ 3.3 Evaluation on Classification
Results. In scenario (1), from Tab. 4, we can observe that the classification results on HWOBC and Oracle-50k datasets are markedly better than those on the OBI125 dataset (surpass by 16.4% and 19.2% for GPT-4o under ‘Yes-or-No’ queries). We believe that this can be attributed to the image quality of OBI in the dataset, which affects the LMM’s perception of character structure. Indeed, the two handprinted OBI datasets, HWOBC and Oracle-50k, are cleaner than the noise-contaminated OBI125 dataset (Sec. C.3). In scenario (2), we find significant performance discrepancies between ‘Yes-or-No’ and ‘How’ queries, which is similar to the other tasks. This highlights the necessity of designing more appropriate textual answer forms and task-specific metrics. Moreover, the Gemini 1.5 series is on par with that of their best competitors (GPT-4v and GPT-4o), which even achieves nearly 90% and 100% accuracy in terms of Acc@1 and Acc@5 metrics respectively. This proves the feasibility of using LMMs for preliminary classification of OBI images. In scenario (3), we select four representative LMMs, i.e., InternVL2-8B, InternVL2-40B, InternVL2-Llama3-76B, and GPT-4v, for evaluation, and obtain three observations from Fig. 3. First, we notice an overall performance improvement compared to the previous experimental setup since the source-mixed test set increases the inter-class differences. Second, the accuracy of these models decreases as the number of classes increases. This shows that current LMMs struggle to perform structural comparisons across a large number of categories, potentially limited by the granularity of their outputs. Third, compared to models with smaller scale, the performance degradation of InternVL2-Llama3-76B and GPT-4v is less pronounced as the number of classification categories increases, indicating that LMMs with larger parameter sizes tend to have better robustness.
@@ 3.4 Evaluation on Retrieval
Results. As presented in Tab. 5, most proprietary LMMs can well retrieve relevant OBI contents from test pools except GLM-4v, which is on average 40.48% worse than the other five in terms of Recall@1. Besides, when expanding the search scope to ten candidates, all proprietary LMMs as well as over 63.6% open-source LMMs perfectly retrieve all relevant OBI contents in OBI-IJDH. However, the retrieval performance fell off a cliff in the OBC306 dataset, which we attribute to the same image quality problems encountered with the classification task. As visualized in Fig. 1, such cropped oracle bone rubbings contain various distortion problems, such as stroke-broken, edge-cracked, spindles, and dense white regions, making them harder to tell apart and showing great necessity to add image enhancement measures before these tasks. Furthermore, models with large number of parameters still perform relatively better in this task and How queries receive more precise search results.
@@ 3.5 Evaluation on Deciphering
Results. For scenario (1), Tab. 6 presents the evaluation results across different character frequencies. We find that the evaluated LMMs achieve on average 0.1905, 0.1898, and 0.1791 on Tier1, Tier2, and Tier3 subsets, respectively, indicating that the deciphering performance on common characters is better than that on rare characters. For scenario (2) and (3), the deciphering results on pictograph are more accurate than ideogram (average BERTScore: 0.2012 > 0.1837), which shows the characteristics of pictographs evolving from drawings and reveals the deciphering mechanism of LMMs, that is, by establishing visual associations between font structure and real objects. For scenario (4), we notice the worst performance (average BERTScore = 0.1636) on the Radical subset compared to other test sets, which means LMMs are impeded in distinguishing and deciphering component-level variants. Furthermore, proprietary LMMs outperform open-source LMMs by an average of 145.9% across all deciphering tasks, suggesting that the open-source LMMs still have a substantial gap to fill for practical application. It is worth noting that GPT-4o, GPT-4v, and Gemini 1.5 Pro have approached or exceeded public-level humans in some scenarios. Surprisingly, InterLM-XComposer2-VL reaches the Top-1 performance among all open-source LMMs (even surpassing InternVL2-Llama3-76B and InternVL2-40B which owns several times of parameters). We suspect that this might be attributed to the specific training datasets (Cai et al., 2024), since the match of data content sometimes outweighs the sheer quantity of parameters (Zhang et al., 2024a). Meanwhile, we observe some highly repetitive outputs (such as ‘The image you’ve provided appears to be a stylized representation of an oracle bone character.’ or other meaningless answers) in the LLaVA-NeXT series and its previous version LLaVA-v1.5, which leads to inferior deciphering results in all scenarios. Fig. 4 provides an example of the deciphering results of GPT-4o and Qwen-VL-Max on four high-frequency oracle bone characters. We also investigate the effect of adding pre-prompt during querying in App. C.4.1. More deciphered results for currently undeciphered OBI are presented in App. C.4.2.

### v2

@@ Abstract
We introduce OBI-Bench, a holistic benchmark crafted to systematically evaluate large multi-modal models (LMMs) on whole-process oracle bone inscriptions (OBI) processing tasks demanding expert-level domain knowledge and deliberate cognition. OBI-Bench includes 5,523 meticulously collected diverse-sourced images, covering five key domain problems: recognition, rejoining, classification, retrieval, and deciphering. These images span centuries of archaeological findings and years of research by front-line scholars, comprising multi-stage font appearances from excavation to synthesis, such as original oracle bone, inked rubbings, oracle bone fragments, cropped single characters, and handprinted characters. Unlike existing benchmarks, OBI-Bench focuses on advanced visual perception and reasoning with OBI-specific knowledge, challenging LMMs to perform tasks akin to those faced by experts. The evaluation of 6 proprietary LMMs as well as 17 open-source LMMs highlights the substantial challenges and demands posed by OBI-Bench. Even the latest versions of GPT-4o, Gemini 1.5 Pro, and Qwen-VL-Max are still far from public-level humans in some fine-grained perception tasks. However, they perform at a level comparable to untrained humans in deciphering tasks, indicating remarkable capabilities in offering new interpretative perspectives and generating creative guesses. We hope OBI-Bench can facilitate the community to develop domain-specific multi-modal foundation models towards ancient language research and delve deeper to discover and enhance these untapped potentials of LMMs.
@@ 1 Introduction
Nowadays, the emergent large multi-modal models (LMMs) have brought opportunities for solving multidisciplinary tasks with powerful visual perception, understanding, and reasoning abilities (Yue et al., 2024; Caffagni et al., 2024; OpenAI, 2024b). Meanwhile, the ability of LMMs remains unclear on fine-grained perception and cognition (i.e., high-level perception), which play significant roles in interpreting ancient texts and their associated tasks on image material pre-processing. The perception at different granularities is strongly associated with a wide range of OBI tasks, such as distinguishing noise from real oracle bone character recognition (Wang et al., 2022), marginal sealability checking in rejoining (Zhang et al., 2021c) and component-level structural decomposition in retrieval (Hu et al., 2024). Henceforth, it is worth evaluating the current abilities of LMMs in fine-grained perception and cognition based on actual needs to relieve extensive human resources and seek feasible solutions for future OBI research.
@@
To systematically evaluate the whole-process perception ability on various visual granularity under diverse OBI tasks, we collect 5,523 OBI images from 11 distinct sources. Due to the lack of publicly available OBI recognition datasets on real oracle bones and OBI rejoining datasets, we propose the original oracle bone recognition (O2BR) dataset and OBI-rejoin dataset as complements. Aligned with existing practices (Yue et al., 2024; Wu et al., 2024), each image in OBI-Bench is equipped with a question alongside a correct answer. Moreover, we craft a spectrum of questions: What question, Yes-or-No question, How question, and Where question, evaluating from coarse-grained perception to finer-grained perception. A total of 23 popular LMMs are selected for evaluation including 6 proprietary LMMs (e.g., GPT-4o, Qwen-VL-Max, and Gemini 1.5 Pro) and 17 open-source LMMs (e.g., LLaVA-Next, InternVL2, and mPLUG-Owl3). We showcase some empirical findings here: 1) LMMs still have much room for improvement in fine-grained OBI recognition tasks such as quantity detection or locating; 2) LMMs are not sensitive enough to local information such as the borders of the fragmented margin to directly meet the requirements of rejoining. Meanwhile, part of LMMs can help OBI scholars narrow down the range of qualified fragments (GPT-4o and Gemini 1.5 Pro achieve an average performance of 77.57% on the Acc@10 metric); 3) LMMs can be applied in classification and retrieval tasks, especially for cleaner handprinted datasets; 4) partial LMMs rival the performance of public-level humans in deciphering tasks and exhibit remarkable visual reasoning ability in interpreting those genuine undeciphered oracle bone characters. The full list of our findings from different OBI tasks is in Sec. 3.
@@ 2.1 General Principles
Covering Multi-stage Font Appearances. To cover the diverse appearances of the oracle bones since their excavation, we collect multi-sourced images for each task, as depicted in Tab. 1. In particular, considering that there are currently no available original oracle bone recognition datasets and rejoining datasets, we proposed the O2BR and OBI-rejoin datasets respectively (See more details in App. B). Due to the specific requirement of domain knowledge in OBI research, three domain experts and one senior OBI scholar were involved in the annotation process of O2BR and OBI-rejoin as well as the sampling process of all source datasets. We include all the manifestations of the entire process of OBI processing from excavation (i.e., original oracle bone) to artificial synthesis (i.e., handprinted or computer-generated). The diverse and multiple sources of images morph the OBI-Bench into a holistic and balanced benchmark to fairly evaluate the competency of LMMs in OBI-related tasks.
@@ 3.1 Evaluation on Recognition
Setup. To evaluate how well LMMs perform under various OBI recognition scenarios, we design four evaluation schemes across two image sources (i.e., original oracle bone from our self-curated O2BR dataset and inked rubbings from the YinQiWenYuandetection (AYNU, 2020) dataset): (1) evaluating via the What query to measure the preciseness of the global perceptual descriptions on OBIs. Given the commonality of contents in sampled test sets, we predefined reference descriptions for each of the two datasets. The max_tokens for all LMMs is set to $200$. The similarity between text embeddings is used as the evaluation metric. (2) evaluating via the Yes-or-No query after specifying the response direction (the words “OBI” appear explicitly in the query), designed to complement the evaluation on coarse-grained perception. The accuracy of answers is adopted as the evaluation criterion. (3) evaluating via the How query to quantify the ability of LMMs in parsing irregularly structured character-level OBI contents. We use mean relative error (MRE) as the metric. (4) evaluating via the Where query to achieve finer-grained perception (i.e., dense oracle bone character locating), designed to more effectively uncover the practicability of LMMs in real OBI processing pipelines. Mean intersection over union (mIoU) is used as the metric. More details are in App. C.2.
@@
Results. In scenario (1), as shown in Tab. 2, among proprietary LMMs, the recently-released GPT-4o reaches the best performance in terms of the relevance of answers on coarse-grained perception, followed by Qwen-VL-Max, which shows rather close results. By achieving over 12.4% on average performance improvement, these models provide more precise answers than the open-source LMMs. Similar improvement can be observed for the same series of LMM (e.g., LLaVA and InternVL2), as the number of parameters of the language backbone increases. In scenario (2), we add keyword (i.e., oracle bone) directly to a Yes-or-No question to further evaluate the global perception of LMMs. We find most LMMs perform nearly 100% correctness, indicating that these models are highly effective in handling queries that involve explicitly directed content. One exception is the GLM-4V series that only achieves an average accuracy of 34.34%. We speculate it is due to the answer preference bias (Wu et al., 2024). A deeper analysis is in App. C.2.2. In scenario (3), we transition to examining the fine-grained visual perceptual of LMMs, i.e., counting the number of oracle bone characters. Our results in Tab. 2 indicate that GLM-4V and InternVL2-Llama3-76B exhibit the best performance within their respective division groups. Additionally, we observe some abnormally large values (>50) or repeated values from Gemini 1.5 Flash, xGen-MM-4B, LLaVA-NeXT-72B, LLaVA-v1.5-13B, and Qwen-VL that result in their relatively large MRE (Zhou et al., 2024). In scenario (4), all evaluated LMMs consistently underperform. Gemini 1.5 Pro significantly outperforms other LMMs (surpasses GPT-4v and GPT-4o two orders of magnitude) (Ueno & Lynn, 2024) with its exclusive coordinate return function. Overall, current LMMs are still not usable for character-level OBI locating and are far from the public-level human.
@@ 3.2 Evaluation on Rejoining
Setup. Given the absence of publicly available rejoining datasets in the OBI domain, we create a dataset consisting of 200 complete OBI pieces across two appearances (i.e., original and inked oracle bone fragments) with 483 rejoinable fragments. More details are in App. B.2. To evaluate the rejoining performance of LMMs on oracle bone fragments with different marginal sealability concerns, we elaborate two evaluation scenarios: (1) evaluating via standard Yes-or-No question that represents the absolute judgments to tentatively explore the rejoining performance; (2) evaluating via How question that outputs rejoinable probabilities for pairs of fragments, allowing for a more accurate and quantitative measurement. Accuracy@k indices are utilized as evaluation metrics.
@@ 3.3 Evaluation on Classification
Results. In scenario (1), from Tab. 4, we can observe that the classification results on HWOBC and Oracle-50k datasets are markedly better than those on the OBI125 dataset (surpass by 16.4% and 19.2% for GPT-4o under ‘Yes-or-No’ queries). We believe that this can be attributed to the image quality of OBI in the dataset, which affects the LMM’s perception of character structure. Indeed, the two handprinted OBI datasets, HWOBC and Oracle-50k, are cleaner than the noise-contaminated OBI125 dataset (Sec. C.3). In scenario (2), we find significant performance discrepancies between ‘Yes-or-No’ and ‘How’ queries, which is similar to the other tasks. This highlights the necessity of designing more appropriate textual answer forms and task-specific metrics. Moreover, the Gemini 1.5 series is on par with that of their best competitors (GPT-4v and GPT-4o), which even achieves nearly 90% and 100% accuracy in terms of Acc@1 and Acc@5 metrics respectively. This proves the feasibility of using LMMs for preliminary classification of OBI images. In scenario (3), we select four representative LMMs, i.e., InternVL2-8B, InternVL2-40B, InternVL2-Llama3-76B, and GPT-4v, for evaluation, and obtain three observations from Fig. 3. First, we notice an overall performance improvement compared to the previous experimental setup since the source-mixed test set increases the inter-class differences. Second, the accuracy of these models decreases as the number of classes increases. This shows that current LMMs struggle to perform structural comparisons across a large number of categories, potentially limited by the granularity of their outputs. Third, compared to models with smaller scales, the performance degradation of InternVL2-Llama3-76B and GPT-4v is less pronounced as the number of classification categories increases, indicating that LMMs with larger parameter sizes tend to have better robustness.
@@ 3.4 Evaluation on Retrieval
Results. As presented in Tab. 5, most proprietary LMMs can well retrieve relevant OBI contents from test pools except GLM-4v, which is on average 40.48% worse than the other five in terms of Recall@1. Besides, when expanding the search scope to ten candidates, all proprietary LMMs as well as over 63.6% open-source LMMs perfectly retrieve all relevant OBI contents in OBI-IJDH. However, the retrieval performance fell off a cliff in the OBC306 dataset, which we attribute to the same image quality problems encountered with the classification task. As visualized in Fig. 1, such cropped oracle bone rubbings contain various distortion problems, such as stroke-broken, edge-cracked, spindles, and dense white regions, making them harder to tell apart and showing great necessity to add image enhancement measures before these tasks. Furthermore, models with a large number of parameters still perform relatively better in this task and How queries receive more precise search results.
@@ 3.5 Evaluation on Deciphering
Results. For scenario (1), Tab. 6 presents the evaluation results across different character frequencies. We find that the evaluated LMMs achieve on average 0.1905, 0.1898, and 0.1791 on Tier1, Tier2, and Tier3 subsets, respectively, indicating that the deciphering performance on common characters is better than that on rare characters. For scenario (2) and (3), the deciphering results on pictograph are more accurate than ideogram (average BERTScore: 0.2012 > 0.1837), which shows the characteristics of pictographs evolving from drawings and reveals the deciphering mechanism of LMMs, that is, by establishing visual associations between font structure and real objects. For scenario (4), we notice the worst performance (average BERTScore = 0.1636) on the Radical subset compared to other test sets, which means LMMs are impeded in distinguishing and deciphering component-level variants. Furthermore, proprietary LMMs outperform open-source LMMs by an average of 145.9% across all deciphering tasks, suggesting that the open-source LMMs still have a substantial gap to fill for practical application. It is worth noting that GPT-4o, GPT-4v, and Gemini 1.5 Pro have approached or exceeded public-level humans in some scenarios. Surprisingly, InterLM-XComposer2-VL reaches the Top-1 performance among all open-source LMMs (even surpassing InternVL2-Llama3-76B and InternVL2-40B which owns several times of parameters). We suspect that this might be attributed to the specific training datasets (Cai et al., 2024) since the match of data content sometimes outweighs the sheer quantity of parameters (Zhang et al., 2024a). Meanwhile, we observe some highly repetitive outputs (such as ‘The image you’ve provided appears to be a stylized representation of an oracle bone character.’ or other meaningless answers) in the LLaVA-NeXT series and its previous version LLaVA-v1.5, which leads to inferior deciphering results in all scenarios. Fig. 4 provides an example of the deciphering results of GPT-4o and Qwen-VL-Max on four high-frequency oracle bone characters. We also investigate the effect of adding pre-prompt during querying in App. C.4.1. More deciphered results for currently undeciphered OBI are presented in App. C.4.2.